OpenStudy (idku):

series representation for tan^(-1)x

OpenStudy (amistre64):

whats the derivative of tan^-1 x ?

OpenStudy (idku):

I am going to start from the elementary geometric series. \(\large\color{black}{ \displaystyle \frac{1}{1-x}=\sum_{n=0}^{\infty}x^n}\) making a substitution of -x instead of x on both sides. \(\large\color{black}{ \displaystyle \frac{1}{1-(-x)}=\sum_{n=0}^{\infty}(-x)^n}\) \(\large\color{black}{ \displaystyle \frac{1}{x+1}=\sum_{n=0}^{\infty}(-1)^nx^n}\) substitution of x^2 instead of x \(\large\color{black}{ \displaystyle \frac{1}{x^2+1}=\sum_{n=0}^{\infty}(-1)^nx^{2n}}\)

OpenStudy (idku):

then I integrate both sides

OpenStudy (idku):

\(\large\color{black}{ \displaystyle \int \frac{1}{x^2+1} dx=\int \sum_{n=0}^{\infty}(-1)^nx^{2n}dx}\)

OpenStudy (idku):

\(\large\color{black}{ \displaystyle \tan^{-1}x=\sum_{n=0}^{\infty}\left[(-1)^n\frac{x^{2n+1}}{2n+1}\right]+C}\)

geerky42 (geerky42):

C shouldn't be included; \(\large\color{black}{ \displaystyle \int \frac{1}{x^2+1} dx=\int \sum_{n=0}^{\infty}(-1)^nx^{2n}dx}\) \(\large\color{black}{ \displaystyle \tan^{-1}x~+C=\sum_{n=0}^{\infty}\left[(-1)^n\frac{x^{2n+1}}{2n+1}\right]+C}\) \(\large\color{black}{ \displaystyle \tan^{-1}x=\sum_{n=0}^{\infty}\left[(-1)^n\frac{x^{2n+1}}{2n+1}\right]}\)

OpenStudy (idku):

but those are not necessarily the same C's on v=both sides

geerky42 (geerky42):

Yeah but C isn't supposed to be there, I just don't know how to explain lol...

OpenStudy (amistre64):

tan^-1 (x) is not a family of functions ...

OpenStudy (amistre64):

you are trying to detemrine the powere series of a specific function

OpenStudy (amistre64):

tan^-1(0) = power series at zero if you want to "solve for C"

OpenStudy (idku):

So no +C ever, in these cases?

OpenStudy (amistre64):

you arent defining a family of functions, so its not pertinent

OpenStudy (idku):

so lets say ln(1-x) \(\large\color{black}{ \displaystyle \ln(1-x)=\sum_{n=0}^{\infty}\left[\frac{x^{n+1}}{n+1}\right]}\)

OpenStudy (amistre64):

int [1/(x^2+1)] dx = tan^-1 (x) +C but C = 0 when we are looking for the value of tan^-1 (x) right?

OpenStudy (idku):

ok, as I understand we don't put the +C here....

OpenStudy (idku):

yes, right...

OpenStudy (amistre64):

the +C is irrelevant, we can either ignore it, or solve for it .. solving for it will show its equal to 0 if you want to insist on keeping it around :)

OpenStudy (idku):

So, if \(\large\color{black}{ \displaystyle f_1(x)=\sum_{n=0}^{\infty}f_2({x,n})}\) then, \(\large\color{black}{ \displaystyle F_1(x)=\sum_{n=0}^{\infty}F_2({x,n})}\) (without +C, for any function in general), such that F`=f

OpenStudy (idku):

it does look confusing, I can elaborate on what I'm asking.

OpenStudy (idku):

\(\large\color{black}{ \displaystyle f_1(x)=\sum_{n=0}^{\infty}f_2({x,n})}\) where: \(\large\color{black}{ \displaystyle f_1(x)}\) some function y=f(x) \(\large\color{black}{ \displaystyle f_2(x,n)}\) equivalent to f(x) power series representation (function of x and n) \(\large\color{black}{ \displaystyle F_1(x)}\) general antideriv. (not a family of functions) for the f(x) \(\large\color{black}{ \displaystyle F_2(x,n)}\) is the antideriv. (not a family of functions) for the power series \(\large\color{black}{ \displaystyle f_2(x,n)}\) (where both the function and antideriv. function are both functions of x and n) then if I wanted a power series to \(\large\color{black}{ \displaystyle F_1(x)}\) (thatis the antideriv. of \(\large\color{black}{ \displaystyle f_1(x)}\) ), then I could integrate both sides, without adding +C's, and get \(\large\color{black}{ \displaystyle F_1(x)=\sum_{n=0}^{\infty}F_2({x,n})}\)

OpenStudy (amistre64):

\[f(x)=a_0+a_1x+a_2x^2+a_3x^3+a_4x^4+...\] because integration is a linear thing \[\int f(x)=\int a_0+\int a_1x+\int a_2x^2+\int a_3x^3+\int a_4x^4+...\] \[F(x)+c_0=a_0x+c_1+\frac12a_1x^2+c_2+\frac13a_2x^3+c_3+\frac14 a_3x^4+c_4+\frac15a_4x^5+c_5+...\] \[F(x)+c_0=(a_0x+\frac12a_1x^2+\frac13a_2x^3+\frac14 a_3x^4+\frac15a_4x^5+...)+k_0\] what is F(x)? well, F(x) + 0 = F(x) so c_0 = 0 \[F(x)=(a_0x+\frac12a_1x^2+\frac13a_2x^3+\frac14 a_3x^4+\frac15a_4x^5+...)+k_0\] now what is F(0)? if F(0) = 3, then k_o = 3 seems fair but k_o is just the constant term of the power series

OpenStudy (idku):

so, no +C :) tnx

OpenStudy (idku):

even if I wrote on my teacher wouldn't deduct, but it is still nice to know.... tnx again

OpenStudy (amistre64):

yep ...

OpenStudy (amistre64):

more succinctly, i would say k_o = F(0), which just represents the constant term of the power series. or k_o = F(a) for some series centered around x=a

OpenStudy (idku):

I am starting to make sense from the math I am doing:D Don't ever or going to be very good at this, but at least I can be decent:) tnx for the help

OpenStudy (amistre64):

good luck :)